# Minimal thermodynamic cost of computing with circuits

### v1

@@ V.1 MMC of different circuit families implementing the same function family
The problem of computing a Boolean function with a circuit family under resource constraints naturally leads to the challenge of optimizing both circuit size and depth complexity. Since a Boolean function can be computed by multiple circuit families, size and depth serve as key measures of computational efficiency. The goal, therefore, is to identify the most efficient circuit family—one that minimizes both size and depth complexity. While size complexity corresponds to space efficiency and depth complexity to time efficiency, mismatch cost complexity serves as a measure of thermodynamic efficiency—the minimal dissipated heat in a circuit. Thus, when comparing two circuit families computing the same Boolean function, mismatch cost is a key metric for evaluating thermodynamic efficiency.
Figure 9: The MMC of two different circuit families implementing the addition function, which takes two integers in the standard $n$-bit representation and returns their sum in the standard $(n+1)$-bit representation.
@@
In Fig. 9, we compare the mismatch cost of RCA and CLA across various input sizes. The results show a linear scaling of mismatch cost with size complexity for both circuits. However, as input size increases, RCA consistently exhibits a lower mismatch cost than CLA, indicating that despite being slower, RCA is thermodynamically more efficient.
@@ V.2 MMC of a circuit family for heterogeneous priors
Fig. 10 illustrates how the mismatch cost of various individual logic gates changes as the prior distribution shifts. Notably, the mismatch cost increases as the prior distribution deviates further from the uniform distribution. Similarly, Fig. 11 shows how the total mismatch cost (MMC) of a circuit composed of AND and OR gates varies as the prior distributions of the two gate types become increasingly heterogeneous.
@@ VI Discussion and Future work
VII Acknowledgement
This work was supported by the U.S. National Science Foundation (NSF) Grant 2221345. We thank Santa Fe Institute for helping to support this research. We thank Harrison Hartle for valuable discussions and insightful feedback during the development of this work.
@@ A.1 Subsystem process
A.2 Mismatch Cost lower bound on EP of a subsystem process
@@
where $\Delta\mathrm{I}=\mathrm{I}_{1}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)-\mathrm{I}_{0}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)$ is the drop in mutual information between the subsystem consisting of $\boldsymbol{x}_{\mu}$ and $\boldsymbol{x}_{\mathrm{pa}(\mu)}$ and the rest of the system.
A.3 Application of subsystem mismatch cost to circuits

### v2

@@ V.1 MMC of different circuit families implementing the same function family
The problem of computing a Boolean function with a circuit family under resource constraints naturally leads to the challenge of optimizing both circuit size and depth complexity. Since a Boolean function can be computed by multiple circuit families, size and depth serve as key measures of computational efficiency. The goal, therefore, is to identify the most efficient circuit family—one that minimizes both size and depth complexity. While size complexity corresponds to space efficiency and depth complexity to time efficiency, mismatch cost complexity serves as a measure of thermodynamic efficiency—the minimal dissipated heat in a circuit. Thus, when comparing two circuit families computing the same Boolean function, MMC is a key metric for evaluating thermodynamic efficiency.
Figure 11: MMC of two different circuit families implementing the addition function, which takes two integers in the standard $n$-bit representation and returns their sum in the standard $(n+1)$-bit representation.
@@
In Fig. 11, we compare MMC of RCA and CLA across various input sizes. The results show a linear scaling of mismatch cost with size complexity for both circuits. However, as input size increases, RCA consistently exhibits a lower MMC than CLA, indicating that despite being slower, RCA is thermodynamically more efficient.
@@ V.2 MMC of a circuit family for heterogeneous priors
Fig. 12 illustrates how the mismatch cost of various individual logic gates changes as the prior distribution shifts. Notably, the mismatch cost increases as the prior distribution deviates further from the uniform distribution. Similarly, Fig. 13 shows how the total MMC of a circuit composed of AND and OR gates varies as the prior distributions of the two gate types become increasingly heterogeneous.
@@ VI Discussion and Future work
VII Acknowledgment
This work was supported by the U.S. National Science Foundation (NSF) Grant 2221345. We thank the Santa Fe Institute for helping to support this research.
@@ A.1 Subsystem process
A.2 MMC lower bound on EP of a subsystem process
@@
where $\Delta\mathrm{I}=\mathrm{I}_{1}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)-\mathrm{I}_{0}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)$ is the drop in mutual information between the subsystem consisting of $\boldsymbol{x}_{\mu}$ and $\boldsymbol{x}_{\mathrm{pa}(\mu)}$ and the rest of the system. Note that Eq. 48 provides an alternative derivation of the mutual information contribution to entropy production in systems with multiple communicating components—such as those involved in feedback control protocols—as previously derived in [26, 27] and [28]. Specifically, it is important to emphasizes that the contribution to entropy production arising from the making and breaking of correlations among components in a non-equilibrium process is captured by MMC.
A.3 Application of subsystem MMC to circuits
@@ Appendix B Proofs
B.3 Circuit as a periodic process
A circuit can be viewed as a periodic process when a counter is integrated into its description. The counter can differentiate one clock cycle from another and sequentially run different layers or gates with each clock cycle. The periodic process, governed by the clock cycle, repeats over and over again, but sequentially implements different gates or layers with each clock cycle.
Let $Z$ denote the random variable representing the state of the counter, taking values $z\in\{0,1,\dots\}$. The joint state of the system is given by $(\boldsymbol{x},z)$, where $\boldsymbol{x}$ is the state of the circuit and $z$ is the state of the counter. We let $\mu$ index the sequence of gates or layers in the circuit. Let $p_{XZ}(\boldsymbol{x},z)$ denote the probability distribution over the joint state, which can be expressed as a conditional distribution of $\boldsymbol{x}$ given the state of the counter:
$p_{X,Z}(\boldsymbol{x},z)=p_{X|Z}(\boldsymbol{x}|z)p_{Z}(z)$ (70)
At the beginning of a circuit run, the counter is initialized to $\mu=0$. With each gate execution, the counter increments by one. Therefore, prior to the execution of gate $\mu$, the counter is in state $z=\mu$. The distribution over circuit states just before the execution of gate $\mu$ is then given by
$p_{X,Z=\mu}(\boldsymbol{x})=p_{X|Z=\mu}(\boldsymbol{x}),$ (71)
indicating that the counter is precisely in state $\mu$. Note that $p_{X|Z=\mu}(\boldsymbol{x})$ corresponds to $p^{\mu}(\boldsymbol{x})$, consistent with the notation used throughout the paper and as summarized in Table 1.
The prior distribution $q_{X,Z}$ represents the full full process that is repeated with each clock cycle of the circuit. Accordingly, the prior over the joint state $(\boldsymbol{x},z)$ can be expressed as
$q_{X,Z}(\boldsymbol{x},z)=q_{X|Z}(\boldsymbol{x}\mid z),q_{Z}(z),$ (72)
and evolves under the periodic dynamics of the circuit to
$\tilde{q}_{X,Z}(\boldsymbol{x},z)=\tilde{q}_{X|Z}(\boldsymbol{x}\mid z),\tilde{q}_{Z}(z),$ (73)
where $\tilde{q}_{X|Z}$ denotes the updated distribution over circuit states following one cycle of evolution.
The prior distribution $q^{\mu}$, associated with the execution of gate $\mu$, is now understood as being conditioned on the counter state. In particular, the prior distribution over circuit states just before the execution of gate $\mu$ (i.e., when the counter is in state $z=\mu$) is given by
$q_{X,Z=\mu}(\boldsymbol{x})=q_{X|Z}(\boldsymbol{x}\mid z),q_{Z=\mu}(z),$ (74)
where $q_{Z=\mu}(z)=\delta_{z,\mu}$ is a delta distribution, enforcing that the counter is precisely in state $\mu$. Equivalently, this can be written more concisely as
$q_{X,Z=\mu}(\boldsymbol{x})=q_{X|Z=\mu}(\boldsymbol{x}).$ (75)
Applying the MMC formula to the execution of gate $\mu$, we obtain
$\mathcal{MC}_{\mu}=D\left(p_{X|Z=\mu}\|q_{X|Z=\mu}\right)-D\left(p_{X|Z=\mu+1}\|\tilde{q}_{X|Z=\mu}\right),$ (76)
which, under the earlier notation $p^{\mu}=p_{X|Z=\mu}$, $q^{\mu}=q_{X|Z=\mu}$, and $\tilde{q}^{\mu}=\tilde{q}_{X|Z=\mu}$, becomes
$\mathcal{MC}_{\mu}=D(p^{\mu}\|q^{\mu})-D(p^{\mu+1}\|\tilde{q}^{\mu}),$ (77)
which coincides with Eq. 23.